NicktheBlade16:

PLZ HELP: Explain how to solve 3x − 4 = 6 using the change of base formula log base b of y equals log y over log b. Include the solution for x in your answer. Round your answer to the nearest thousandth.

Alexis1415:

So is this one just about explaining it or are there choices?

NicktheBlade16:

wants you to explain it

Alexis1415:

So if it's explaining, then the answer would be... x = 5.631 because, \[3^(x-4) = 6\] Then, Make sure you log aka use the logarithm, base 3 of both sides. \[3^(x − 4) = \log₃ 6\] \[x − 4 = \log₃ 6\] After that make sure you use the correct change of the base formula (sometimes you can switch it up and make a mistake by using the wrong number). \[x − 4 = \log 6 / \log 3\] And finally we must solve for x. \[x - 4 = \log 6 / \log 3\] Which gives us our answer: 5.631 I hope this helps!

Alexis1415:

(the parenthesis in the problem are supposed to be exponents but it corrected to normal formula)

NicktheBlade16:

Ok, i dont understand it but im sure you are correct

Alexis1415:

Basically the log parts are the inverse function to the exponents so like for example 1000=10 and the base is 10 of 1000 is 3 or like how i would write/type log10 (1000) = 3. does that make a bit more sense?

NicktheBlade16:

actually. yes it does. Thank you again. :)

Alexis1415:

@nicktheblade16 wrote:
actually. yes it does. Thank you again. :)
ofc